OpenStudy (anonymous):

Numbers inside a parenthesis is a fraction, the / being the division bar. Solve for x: (1/x+6)-(4/x)=(2x+6/x^2+6x)

hero (hero):

Is it: \(\large\frac{1}{x+6} - \frac{4}{x} = \frac{2x+6}{x^2 + 6x}\)

OpenStudy (anonymous):

yeah that's it. sorry, i'm brand new to this site so i don't know how to make the equation look like that lol

hero (hero):

You don't need to know how to do that just yet, however, you should at least know how to write fractions linearly

hero (hero):

1/(x+6) - 4/x = (2x+6)/(x^2+6x) How you place the parentheses are important

hero (hero):

I will post a solution soon

OpenStudy (anonymous):

ok thank you

hero (hero):

0. Given \(\frac{1}{x+6} - \frac{4}{x} = \frac{2x+6}{x^2 + 6x}\) 1. Multiply 1/(x+6) by x/x: \(\frac{x}{x^2+6x} - \frac{4}{x} = \frac{2x+6}{x^2 + 6x}\) 2. Subtract x/(x^2 + 6x) from both sides: \( - \frac{4}{x} = \frac{2x+6}{x^2 + 6x} - \frac{x}{x^2+6x}\) 3. Combine the right side: \( - \frac{4}{x} = \frac{2x+6 - x}{x^2 + 6x}\) 4. Simplify the right side: \( - \frac{4}{x} = \frac{x+6}{x^2 + 6x}\) 5. Factor the denominator of the right side fraction: \( - \frac{4}{x} = \frac{x+6}{x(x + 6)}\) 6. Cancel the factor of 1 (x+6)/(x+6): \( - \frac{4}{x} = \frac{1}{x}\) 7. Cross mulitply: \(\small-4x = x\) 8. Add 4x to both sides: \(\small4x + x = 0\) 9. Combine the left side: \(\small5x = 0\) 10. Divide both sides by five: \(\small x = \frac{0}{5}\) 11. Solution \(\small x \ne 0\), therefore no solution.

OpenStudy (anonymous):

Thanks! That helps a lot.

hero (hero):

Yeah, I double checked \(x \ne 0\), so no solution.